OpenStudy (anonymous):

If IAxBI = IA.B|, then the angle between A and B is supposed to be 9O°, but how ? If I do ab sin ☆ = ab cos ☆, then I get ☆ as 45° What am I doing wrong?

OpenStudy (vincent-lyon.fr):

You probably mean IAxBI = IAI.IBI

OpenStudy (irishboy123):

\[|\vec A \times \vec B| = |\vec A|||\vec B|\sin \theta| \] \[|\vec A \bullet \vec B| = |\vec A|||\vec B|\cos \theta| \] if \[|\vec A \times \vec B| =|\vec A \bullet \vec B| \ then\] \[\theta = \frac{\pi}{4}\] ie the original proposition is wrong. if

OpenStudy (irishboy123):

if \[\theta = \frac{\pi}{2} \ then\] \[| \vec A \times \vec B | = |\vec A||\vec B|\]

OpenStudy (irishboy123):

but if \[\theta = \frac{\pi}{2} \ then\] \[\vec A \bullet \vec B = 0, \ so \ |\vec A \bullet \vec B| = 0\]

OpenStudy (anonymous):

If you take theta as 90 then you get ab sin 90 = ab cos 90 , ie ab(1) = ab (0), which are not equal. How do I prove for theta equals 90?

OpenStudy (irishboy123):

"then you get ab sin 90 = ab cos 90" no, you don't, if you apply the formulae *correctly*. read above in short, if theta = 90, then the dot prod A.B is zero and the cross prod A x B is |A||B|. if this is not helping, you might be better posting a source question or an extract from your notes or the basis of the original incorrect proposition.

OpenStudy (thomas5267):

Just a rephrase of what IrishBoy123 said. \[ \begin{align*} \|\vec{A}\|\|\vec{B}\|\sin(\theta)&=\|\vec{A}\|\|\vec{B}\|\cos(\theta)\\ \sin(\theta)&=\cos(\theta)\\ \tan(\theta)&=1\\ \theta&=\frac{\pi}{4} \text{ or }\frac{5\pi}{4},\,0\leq\theta<2\pi \end{align*} \] If the angle between \(\vec{A}\) and \(\vec{B}\) is 90°, then dot product is zero and cross product is \(\|\vec{A}\|\|\vec{B}\|\), so the statement "If \(\|\vec{A}\times\vec{B}\|=\|\vec{A}\cdot\vec{B}\|\), then the angle between A and B is supposed to be 90° " is wrong.

OpenStudy (anonymous):

Ok I've got it! Thank you @Irishboy123 , @thomas5267 and @Vincent-Lyon.Fr :)